# TypedThinker: Typed Thinking Improves Large Language Model Reasoning

### v1

@@ 3.2 Optimize Implicit Policy for Reasoning Type Selection and Following
We use a self-training framework to optimize the meta-thinker $\pi_{\theta}$ and the reasoner while updating the explicit memory with the collected experience. The pipeline is demonstrated in Figure 3. The green lines represent the parametric optimization process, while the blue line represents the non-parameter update.
@@
Diversify Reasoning Experiences with Types To inspire LLMs’ knowledge of solving problems with different reasoning types, the definition (Table 8) and manually-written few-shot examples with detailed reasoning paths (Table 11) are used for prompting solutions for each reasoning type. For each problem in the training set, we use a temperature of 1 to sample 10 solutions per reasoning type. These solutions are then filtered by the correctness of the final answers. To guarantee that these solutions belong to their reasoning type, we apply a reverse check on the remaining solutions. For the experience $(\bm{x},sol,y)$ of the reasoning type $f_{k}$, we prompt the model to predict its reasoning type $\hat{f}_{k}$. If $f_{k}=\hat{f}_{k}$, we think this experience indicates the methodology of this reasoning type and keep it. Otherwise, it will be removed. Finally, we get an experience dataset $\overline{D}$ with multiple reasoning types. The experiences are grouped by their reasoning type and are stored in the explicit memory $M$.
@@
To conclude, TypedThinker first uses the meta-thinker to implicitly identify the reasoning types $f_{k}$ needed to solve the problem, then retrieves relevant demonstrations $d_{k}$ from the explicit memory, and finally employs the fine-tuned reasoner $M_{\phi}$ to execute the specific reasoning type $f_{k}$ on the query with $d_{k}$ to generate the response based on the best reasoning type or a weighted reasoning vote over the effective reasoning set.
@@ 4 Experiments
In this section, we first introduce our experiment settings and investigate the influence of different reasoning types on the self-training data. We then evaluate our TypedThinker performances on four benchmarks. Further analyses are provided to enhance the understanding of TypedThinker.
@@ 4.1 Experiment Setup
We investigate two open-source LLMs Mistral 7B instruct (Jiang et al., 2023) and LLaMA3 8B instruct (Touvron et al., 2023) on two logical benchmarks (LogiQA, BBH) and two mathematics benchmarks (GSM8K and MATH). For each LLM, we set up the following baselines: (i) Few-shot baseline with 3 in-context examples. We use the few-shot examples provided in  Suzgun et al. (2022) for BBH, and text-based few-shot examples in  Toshniwal et al. (2024) for GSM8k and MATH since we do not consider the code interpreter in this paper. We also manually write few-shot examples for LogiQA, (ii) CoT Selection: Select the best reasoning type by prompting. We let the LLM identify the best reasoning type and then apply the selected type to the problem. (iii) Zero-shot Mixture of Reasoning (MoR): apply all possible reasoning types and use the majority vote to get the final answer  [^3] For answers with the same votes, we choose the first one in alphabetical order.. The LLM is instructed with the definition and demonstration in Table 8. (iv) Few-shot MoR: Similar to the zero-shot MoR except for each reasoning type, 3 few-shot examples are provided in the prompt. (v) TypedThinker: use the most effective reasoning type $f^{*}$. The +SC baselines indicate the majority vote over 5 responses.
The temperature is set to 0.7 for all baselines as suggested by  Wang et al. (2022). The maximum output length is set to 1000 tokens. We use SentenceTransformer [^4] https://www.sbert.net/(Reimers and Gurevych, 2019) to retrieve top-3 similar experiences and the threshold is set to $\delta=0.5$. On logical benchmarks, we compare the response with the ground truth based on the exact match. On the mathematics benchmarks, we follow  Toshniwal et al. (2024) to get the accuracy on mathematics benchmarks.
@@ 4.2 How do reasoning types encourage divergent thinking during generation?
Table 1: TypedThinker achieves the best performance in both single response and the majority vote setting on two logical benchmarks and two math benchmarks. @5 indicates the result is based on the majority vote over 5 responses. +SC indicates the self-consistency method. MoR indicates the Mixture of Reasoning, which employs all reasoning types (including an empty type) and votes for the final output. Avg. indicates the average accuracy over four benchmarks.
@@
Few-shot 0.485 0.346 0.369 0.074 0.318 0.566 0.318 0.472 0.105 0.365
+ SC @5 0.532 0.441 0.444 0.136 0.388 0.573 0.364 0.457 0.134 0.382
CoT Selection 0.474 0.361 0.372 0.095 0.325 0.552 0.370 0.353 0.094 0.342
+ SC @5 0.503 0.429 0.466 0.132 0.382 0.583 0.424 0.460 0.134 0.400
Zero-shot MoR @5 0.528 0.414 0.313 0.108 0.341 0.558 0.448 0.261 0.075 0.335
Few-shot MoR @5 0.509 0.456 0.460 0.127 0.388 0.568 0.428 0.470 0.133 0.400
TypedThinker 0.554 0.423 0.386 0.092 0.364 0.550 0.533 0.535 0.193 0.453
+ SC @5 0.570 0.469 0.500 0.149 0.422 0.620 0.591 0.723 0.263 0.549
@@ 4.3 What kinds of benefits can TypedThinker bring?
Appropriate reasoning types improve the reasoning performance. The main difference between Fewshot and CoT Selection without the majority vote is the reasoning type selection. For CoT Selection, the model is first prompted to predict a reasoning type and then apply it, while the Fewshot baseline directly solves the problem. However, we find that the CoT Selection struggles with the reasoning type selection. Given the option to choose from four reasoning types or none, it chooses none over 60% of the time. The rest of the time, it selects more than 50% deductive, while only 34% of them can be effectively solved by deductive reasoning during the sampling. The mismatch in reasoning types results in poor performance. Facilitating with a trained meta-thinker, TypedThinker is more accurate in selecting the reasoning type, which helps it improve performance under the single response setting.
@@
Experience of how to conduct a specific type of reasoning is important. The performance difference between zero-shot and few-shot MoR illustrates the impact of the reasoning demonstration. When prompted solely with the definition, LLMs struggle to understand how to apply the reasoning type to specific problems. It can be improved by human-written few-shot examples in few-shot MoR. However, it still falls behind the non-parametric retrieval and the parametric reasoner in TypedThinker, both of which enhance the capability of conducting a specific reasoning type.
@@ 4.4 What contributes to TypedThinker’s effectiveness?
Ablation study The ablation studies are conducted on three key components in TypedThinker. Each time one component is removed. It includes (i) w/o Fine-tuned Reasoner: it is replaced with the base LLM (ii) w/o Meta-thinker: it is replaced with a CoT selection (iii) w/o Memory: the explicit memory is replaced with the human-written few-shot examples of each reasoning type. In Table 2, we can find the meta-thinker is the most important module for the math benchmarks, while the explicit memory is more effective on two logical benchmarks. The fine-tuned reasoner also contributes a lot to the performance improvement. We also observe that explicit memory does not always bring benefits: the performance on MATH even increases when we remove it. We find that the retrieved examples usually have a similar context but different numbers. The math calculation in the retrieved chain-of-thoughts solutions will mislead the reasoner. This is consistent with the observations of  Toshniwal et al. (2024) that the solutions with masked computations are more beneficial to the math problems. For logical problems, there are fewer calculations and the retrieved solutions focus more on the reasoning process.
@@
Optimal reasoning type v.s. weighted vote on the effective set In the main experiment, we use the optimal reasoning type $f^{*}$ which has the highest effectiveness score for reasoning. As discussed in Section 3.1, we can also use a majority vote on the effective set $F$ with the effectiveness score as the coefficient. Specifically, if one solution is based on a reasoning type with a higher effectiveness score, its vote gets a larger weight. The results are shown in Figure 4. We can see that the weighted vote can balance different reasoning types on LogiQA and GSM8k for the Mistral-7B-based model. However, on the other two benchmarks, the TypedThinker + SC @5 has a better performance. It indicates that accurate selection is more important if one or two reasoning types dominate the benchmark. For example, as we have shown in Figure 1, there are a lot of problems that can only be solved by inductive reasoning, indicating the other types will mislead the final answer. In such cases, the self-consistency of inductive reasoning is more powerful than the weighted vote. However, when we have a more accurate meta-thinker that can identify the correct reasoning type and a more powerful reasoner that can follow the specific reasoning type, for example, models initialized by LLaMA3, the advantage of TypedThinker + SC is more obvious.
@@
CoT Selection 0.415 0.369
Zero-shot MoR 0.415 0.357
Few-shot MoR 0.432 0.398
TypedThinker 0.452 0.403
@@ 4.5 Can TypedThinker be applied to new domains or new LLMs without finetuning?
It is essential to evaluate the generalization capability of our TypedThinker framework. We assess it from two aspects: (i) TypedThinker’s performance on new domains; and (ii) other LLMs’ performance after facilitated with our finetuned meta-thinker and the explicit memory.
@@
TypedThinker generalizes well to the unseen domain. We use a new propositional logic benchmark Contexthub (Hua et al., 2024) for evaluation. It is a recently released dataset, which has never been seen by Mistral and LLaMA3 models during the pre-training. It contains problems from 12 categories with 4 levels of difficulty. We select the difficulty of level 4 to test the complex logic reasoning capabilities. We use the experiences collected from LogiQA as the explicit memory. The meta-thinker and the reasoner are fine-tuned on four training benchmarks. The results in Table 4 show that TypedThinker outperforms other baselines on this unseen domain as well, indicating that it can generalize well to new domains. One interesting thing is Mistral 7B baselines significantly outperform LLaMA3 8B on this benchmark and its superior capabilities make it benefit more from our TypedThinker.
@@
Facilitating LLMs with TypedThinker makes them more powerful. Our TypedThinker framework is orthogonal to the backbone LLMs and can be adapted to new LLMs. There are two ways to use a new LLM in the TypedThinker framework: one is to conduct the self-training process, like the two LLMs used in our main experiments (Mistral 7B and LLaMA3 8B); the other is to use our fine-tuned meta-thinker for reasoning type selection and the explicit memory for retrieval while using the new LLM as the reasoner without finetuning. The first way can make the LLM more powerful (as shown in the performance comparison between TypedThinker and TypedThinker w/o Finetuned Reasoner in Table 2), but the latter one is more flexible. Here we use the second way to evaluate the direct transferability to new LLMs. We choose one of the most powerful LLMs GPT-4o and one math-specific 7B model MetaMath (Yu et al., 2023). MetaMath is a Mistral-7B-based model trained with more than 400k synthesized math data distilled from GPT-3.5-Turbo. We randomly sample 100 examples from each benchmark for GPT-4o. For MetaMath, we use the whole test set. The results are shown in Table 6 and Table 6. Compared with Mistral 7B in Table 1, the high-quality and large scale of synthesized data from GPT-3.5-Turbo enhances MetaMath’s capabilities in math problems. TypedThinker can further improve its performance by reasoning type selection and explicit memory. Meanwhile, although the superior performance of GPT-4o on two math datasets leaves little space for improvement, the results on logic benchmarks (LogiQA and BBH) demonstrate that the meta-thinker trained with the small 7B model also enhances its performance. These findings confirm that our approach is not only effective in improving smaller LLMs but also transferable to larger models, further validating the generalization capability of TypedThinker.
@@
Table 5: GPT-4o’s performance is improved with our meta-reasoner. We use the finetuned Mistral 7B meta-reasoner to predict the reasoning type and replace the fine-tuned reasoner with GPT-4o.
LogicQA BBH GSM8k MATH
@@ 4.6 Case study
Here is one example of TypedThinker on the LogiQA benchmark in Table 7. This problem states a phenomenon that a higher altitude leads to a lower atmospheric pressure. Based on this observation, it is easy for humans to use inductive reasoning and get a general conclusion about the inverse cause-and-effect relationship. It is also natural for humans to use analogical reasoning and find the most similar options. The meta-thinker gives the highest effectiveness score for inductive reasoning, which is then chosen as the optimal reasoning type $f^{*}=\text{inductive}$. Effectiveness scores are all larger than 0, so the effective set is all reasoning types. The reasoner gets the correct answer for deductive and inductive reasoning while doing wrong on the other reasoning types. If we use the majority vote over five answers, (A) and (C) will have the same votes, indicating that there is a 50% chance to be correct  [^5] In our implementation, answers with the same votes are ranked based on their alphabetical order, so (A) will be chosen in this case.. However, with the effectiveness score predicted by the meta-thinker, TypedThinker can get the correct answer either by applying the optimal reasoning type or using the weighted vote on the four answers. Besides, without a specific reasoning type (which is ‘Empty’), the model cannot arrive at the correct answer. This shows the limitation of the common few-shot baselines. It shows that TypedThinker improves the reasoning performance by the introduction of diverse reasoning types and the capability of selecting the appropriate type to apply
Table 7: One example from LogiQA. The correct answer and the reasoning type with the highest effectiveness score are underlined. MoR is the few-shot MoR baselines, which use the majority votes among all reasoning types.
Problem The higher the altitude, the smaller the atmospheric pressure. Because the altitude of Lanzhou is higher than that of Tianjin, the atmospheric pressure of Lanzhou is lower than that of Tianjin. Which of the following reasoning is most similar?
(A) In a highly competitive market, the better the product quality and the more advertising investment, the greater the product sales. Company A invests more money in advertising than Company B. So company A sells more products than company B
(B) The older a person is, the more mature he becomes. Lao Zhang is older than his son, so Lao Zhang is more mature than his son
(C) The older a tree is, the more rings it has. The age of the locust tree in Lao Zhang’s yard is older than that of Lao Li’s family, so the locust tree of Lao Zhang’s family has more rings than Lao Li’s
(D) The greater the vocabulary of a language, the more difficult it is to learn. English is harder to learn than Italian, so English has a larger vocabulary than Italian

### v2

@@ 3.2 Optimize Implicit Policy for Reasoning Type Selection and Exploiting
We optimize the meta-thinker $\pi_{\theta}$ and the reasoner while updating the explicit collection of demonstration with the collected experience. The pipeline is demonstrated in Figure 3. The green lines represent the parametric optimization process, while the blue line represents the non-parameter update.
@@
Diversify Reasoning Experiences with Types To inspire LLMs’ knowledge of solving problems with different reasoning types, the definition (Table 8) and manually-written few-shot examples with detailed reasoning paths (Table 16) are used for prompting solutions for each reasoning type. For each problem in the training set, we use a temperature of 1 to sample 10 solutions per reasoning type. These solutions are then filtered by the correctness of the final answers. To guarantee that these solutions belong to their reasoning type, we apply a reverse check on the remaining solutions. For the experience $(\bm{x},sol,y)$ of the reasoning type $f_{k}$, we prompt the model to predict its reasoning type $\hat{f}_{k}$. If $f_{k}=\hat{f}_{k}$, we think this experience indicates the methodology of this reasoning type and keep it. Otherwise, it will be removed. Finally, we collect an experience dataset $\overline{D}$ with multiple types of reasoning. The experiences are grouped by their reasoning type and are stored in the explicit collection of demonstration $M$.
@@ 4.1 Experiment Setup
We investigate two open-source LLMs Mistral 7B instruct (Jiang et al., 2023) and LLaMA3 8B instruct (Touvron et al., 2023) and Qwen 2-7B-Instruct (Bai et al., 2023) on two logical benchmarks (LogiQA, BBH) and two mathematics benchmarks (GSM8K and MATH). For each LLM, we set up the following baselines: (i) Few-shot baseline with 3 in-context examples. We use the few-shot examples provided in  Suzgun et al. (2022) for BBH, and text-based few-shot examples in  Toshniwal et al. (2024) for GSM8k and MATH since we do not consider the code interpreter in this paper. We also manually write few-shot examples for LogiQA, (ii) CoT Selection: Select the best reasoning type by prompting. We let the LLM identify the best reasoning type and then apply the selected type to the problem. (iii) Self-Discover (Zhou et al., 2024) generates a task-level reasoning structure by prompting LLMs to select relevant modules from a list of seed modules and adapt the selected module to task-specific descriptions. We follow their official implementation  [^3] https://github.com/kailashsp/SELF-DISCOVER and use the backbone LLMs to generate one reasoning structure from an exemplar training instance of each task. This reasoning structure is then applied to all instances in this task. (iv) Zero-shot Mixture of Reasoning (MoR): apply all possible reasoning types and use the majority vote to get the final answer  [^4] For answers with the same votes, we choose the first one in alphabetical order.. The LLM is instructed with the definition and demonstration in Table 8. (v) Few-shot MoR: Similar to the zero-shot MoR except for each reasoning type, 3 few-shot examples are provided in the prompt. (vi) TypedThinker: use the most effective reasoning type $f^{*}$. The +SC baselines indicate the majority vote over 5 responses.
The temperature is set to 0.7 for all baselines as suggested by  Wang et al. (2022). The maximum output length is set to 1000 tokens. We use SentenceTransformer [^5] https://www.sbert.net/(Reimers and Gurevych, 2019) to retrieve top-3 similar experiences and the threshold is set to $\delta=0.5$. We use accuracy as the measurement of task performance. The model response is compared with the ground truth based on the exact match for logic problems. The script in  Toshniwal et al. (2024) calculates the mathematical equivalent of mathematics benchmarks. The training details are put in Appendix A.3.
@@ 4.2 How do reasoning types encourage divergent thinking during generation?
Table 1: TypedThinker achieves the best performance in both single response and the majority vote setting on two logical benchmarks and two math benchmarks. @5 indicates the result is based on the majority vote over 5 responses. +SC indicates the self-consistency method. MoR indicates the Mixture of Reasoning, which employs all reasoning types (including an empty type) and votes for the final output. Avg. indicates the average accuracy over four benchmarks. Qwen’s results are put in Table 11.
@@
Few-shot 0.485 0.346 0.369 0.074 0.318 0.581 0.359 0.581 0.193 0.428
+ SC @5 0.532 0.441 0.444 0.136 0.388 0.579 0.391 0.769 0.250 0.497
CoT Selection 0.474 0.361 0.372 0.095 0.325 0.564 0.392 0.556 0.181 0.423
+ SC @5 0.503 0.429 0.466 0.132 0.382 0.562 0.426 0.785 0.222 0.499
Self-Discover 0.386 0.340 0.141 0.056 0.231 0.493 0.425 0.587 0.200 0.426
+ SC @ 5 0.476 0.391 0.208 0.082 0.289 0.540 0.543 0.701 0.278 0.516
Zero-shot MoR @5 0.528 0.414 0.313 0.108 0.341 0.556 0.463 0.666 0.189 0.468
Few-shot MoR @5 0.509 0.456 0.460 0.127 0.388 0.599 0.543 0.585 0.195 0.481
TypedThinker 0.554 0.423 0.386 0.092 0.364 0.599 0.543 0.585 0.195 0.481
+ SC @5 0.570 0.469 0.500 0.149 0.422 0.637 0.591 0.753 0.267 0.562
@@ 4.3 What kinds of benefits can TypedThinker bring?
Appropriate reasoning types improve the reasoning performance. The main difference between Fewshot and CoT Selection without the majority vote is the reasoning type selection. For CoT Selection, the model is first prompted to predict a reasoning type and then apply it, while the Fewshot baseline directly solves the problem. However, we find that the CoT Selection struggles with the reasoning type selection. Given the option to choose from four reasoning types or none, it chooses none over 60% of the time. The rest of the time, it selects more than 50% deductive, while only 34% of them can be effectively solved by deductive reasoning during the sampling. The mismatch in reasoning types results in poor performance. Facilitating with a trained meta-thinker, TypedThinker is more accurate in selecting the reasoning type, which helps it improve performance under the single response setting. Self-Discover, which uses a shared reasoning structure for all instances of the task, performs poorly, especially for the weaker model Mistral 7B. This may be due to the difficulty these models face in identifying a reusable shared high-level reasoning structure for diverse instances, especially in datasets like LogiQA, where reasoning structures are highly varied.
@@
Experience of how to conduct a specific type of reasoning is important. The performance difference between zero-shot and few-shot MoR illustrates the impact of the reasoning demonstration. When prompted solely with the definition, LLMs struggle to understand how to apply the reasoning type to specific problems. It can be improved by human-written few-shot examples in few-shot MoR. However, it still falls behind the non-parametric retrieval and the parametric reasoner in TypedThinker, both of which enhance the capability of conducting a specific reasoning type. Additionally, poor performance in Self-Discover also indicates that, without demonstration, the complex reasoning structures will introduce excessive complexity, confusing the models.
@@ 4.4 What contributes to TypedThinker’s effectiveness?
Ablation study The ablation studies are conducted on three key components in TypedThinker. Each time one component is removed. It includes (i) w/o Fine-tuned Reasoner: it is replaced with the base LLM (ii) w/o Meta-thinker: it is replaced with a CoT selection (iii) w/o collection of demonstration: the explicit collection is replaced with the human-written few-shot examples of each reasoning type. In Table 2, we can find the meta-thinker is the most important module for the math benchmarks, while the explicit collection is more effective on two logical benchmarks. The fine-tuned reasoner also contributes a lot to the performance improvement. We also observe that explicit collection does not always bring benefits: the performance on MATH even increases when we remove it. We find that the retrieved examples usually have a similar context but different numbers. The math calculation in the retrieved chain-of-thoughts solutions will mislead the reasoner. This is consistent with the observations of  Toshniwal et al. (2024) that the solutions with masked computations are more beneficial to the math problems. For logical problems, there are fewer calculations and the retrieved solutions focus more on the reasoning process.
@@
CoT Selection 0.415 0.390
Zero-shot MoR 0.415 0.403
Few-shot MoR 0.432 0.390
Self-Discover 0.332 0.365
TypedThinker 0.452 0.423
Optimal reasoning type v.s. weighted vote on the effective set In the main experiment, we use the optimal reasoning type $f^{*}$ which has the highest effectiveness score for reasoning. As discussed in Section 3.1, we can also use a majority vote on the effective set $F$ with the effectiveness score as the coefficient. Specifically, if one solution is based on a reasoning type with a higher effectiveness score, its vote gets a larger weight. The results are shown in Table 4. We can see that the weighted vote can balance different reasoning types on LogiQA and GSM8k for the Mistral-7B-based model. However, on the other two benchmarks, the TypedThinker + SC @5 has a better performance. It indicates that accurate selection is more important if one or two reasoning types dominate the benchmark. For example, as we have shown in Figure 1, there are a lot of problems that can only be solved by inductive reasoning, indicating the other types will mislead the final answer. In such cases, the self-consistency of inductive reasoning is more powerful than the weighted vote. However, when we have a more accurate meta-thinker that can identify the correct reasoning type and a more powerful reasoner that can follow the specific reasoning type, for example, models initialized by LLaMA3, the advantage of TypedThinker + SC is more obvious.
@@ 4.5 Can TypedThinker be applied to new domains or new LLMs without finetuning?
It is essential to evaluate the generalization capability of TypedThinker. We assess it from two aspects: (i) TypedThinker’s performance on new domains; and (ii) other LLMs’ performance after facilitated with our finetuned meta-thinker and the explicit collection of demonstration.
@@
TypedThinker generalizes well to the unseen domain. We use a new propositional logic benchmark Contexthub (Hua et al., 2024) for evaluation. It is a recently released dataset, which has never been seen by Mistral and LLaMA3 models during the pre-training. It contains problems from 12 categories with 4 levels of difficulty. We select the difficulty of level 4 to test the complex logic reasoning capabilities. We use the experiences collected from LogiQA as the explicit collection of demonstration. The meta-thinker and the reasoner are fine-tuned on four training benchmarks. The results in Table 4 show that TypedThinker outperforms other baselines on this unseen domain as well, indicating that it can generalize well to new domains. One interesting thing is Mistral 7B baselines significantly outperform LLaMA3 8B on this benchmark and its superior capabilities make it benefit more from our TypedThinker.
@@
Facilitating LLMs with TypedThinker makes them more powerful. Our TypedThinker framework is orthogonal to the backbone LLMs and can be adapted to new LLMs. There are two ways to use a new LLM in the TypedThinker framework: one is to conduct the self-training process, like the two LLMs used in our main experiments (Mistral 7B and LLaMA3 8B); the other is to use our fine-tuned meta-thinker for reasoning type selection and the explicit collection of demonstration for retrieval while using the new LLM as the reasoner without finetuning. The first way can make the LLM more powerful (as shown in the performance comparison between TypedThinker and TypedThinker w/o Finetuned Reasoner in Table 2), but the latter one is more flexible. Here we use the second way to evaluate the direct transferability to new LLMs. We choose one of the most powerful LLMs GPT-4o and one math-specific 7B model MetaMath (Yu et al., 2023). MetaMath is a Mistral-7B-based model trained with more than 400k synthesized math data distilled from GPT-3.5-Turbo. We randomly sample 100 examples from each benchmark for GPT-4o. For MetaMath, we use the whole test set. The results are shown in Table 6 and Table 6. Compared with Mistral 7B in Table 1, the high-quality and large scale of synthesized data from GPT-3.5-Turbo enhances MetaMath’s capabilities in math problems. TypedThinker can further improve its performance by reasoning type selection and explicit collection. Meanwhile, although the superior performance of GPT-4o on two math datasets leaves little space for improvement, the results on logic benchmarks (LogiQA and BBH) demonstrate that the meta-thinker trained with the small 7B model also enhances its performance. These findings confirm that our approach is not only effective in improving smaller LLMs but also transferable to larger models, further validating the generalization capability of TypedThinker.
@@
Table 5: GPT-4o’s performance is improved with our meta-reasoner. We use the finetuned Mistral 7B meta-thinker to predict the reasoning type.
LogiQA BBH GSM8k MATH
@@ 4.6 Case study
Here is one example of TypedThinker on the LogiQA benchmark in Table 7. This problem states a phenomenon that a higher altitude leads to a lower atmospheric pressure. Based on this observation, it is easy for humans to use inductive reasoning and get a general conclusion about the inverse cause-and-effect relationship. It is also natural for humans to use analogical reasoning and find the most similar options. The meta-thinker gives the highest effectiveness score for inductive reasoning, which is then chosen as the optimal reasoning type $f^{*}=\text{inductive}$. Effectiveness scores are all larger than 0, so the effective set is all reasoning types. The reasoner gets the correct answer for deductive and inductive reasoning while doing wrong on the other reasoning types. If we use the majority vote over five answers, (A) and (C) will have the same votes, indicating that there is a 50% chance to be correct  [^6] In our implementation, answers with the same votes are ranked based on their alphabetical order, so (A) will be chosen in this case.. However, with the effectiveness score predicted by the meta-thinker, TypedThinker can get the correct answer either by applying the optimal reasoning type or using the weighted vote on the four answers. Besides, without a specific reasoning type (which is ‘Empty’), the model cannot arrive at the correct answer. This shows the limitation of the common few-shot baselines. It shows that TypedThinker improves the reasoning performance by the introduction of diverse reasoning types and the capability of selecting the appropriate type to apply.
Table 7: One example from LogiQA. The correct answer and the reasoning type with the highest effectiveness score are underlined. MoR is the few-shot MoR baselines, which use the majority votes among reasoning types.
Problem The higher the altitude, the smaller the atmospheric pressure. Because the altitude of Place A is higher than that of Place B, the atmospheric pressure of Place A is lower than that of Place B. Which of the following examples shows the same pattern of reasoning?
(A) In a highly competitive market, the better the product quality and the more advertising investment, the greater the product sales. Company A invests more money in advertising than Company B. So Company A sells more products than Company B.
(B) The older a person is, the more mature he becomes. Person A is older than his son, so Person B is more mature than his son.
(C) The older a tree is, the more rings it has. The age of the locust tree in A’s yard is older than that of B’s family, so the locust tree of A’s family has more rings than B’s.
(D) The greater the vocabulary of a language, the more difficult it is to learn. English is harder to learn than Italian, so English has a larger vocabulary than Italian.